When sarah rowed down black river with the current , she took one hour to go four miles. When she rowed back the same distance, at the same rowing speed, but against the current ,her trip required two hours . what is the speed, in miles per hour, of the current in black river ?
Answer by checkley77(12844)

D=RT
4=R*1
R=4 MPH. ROWING WITH THE CURRENT.
4=R*2
R=4/2
R=2 MPH. ROWING AGAINST THE CURRENT.
(4+2)/2=6/2=3 MPH. IS THE ROWING SPEED IN STILL WATER.
PROOF:
3+1=4
3-1=2